Miss Muffet insists on consuming 2 units of whey per 1 unit of curds. The price of curds is $5 and the price of whey is $3, then if Miss Muffet’s income is m. Calculate her demand for curds will be

PLEASE SHOW ALL WORK
Answer #1

price of curd = 5, price of whey = 3

income = m

Let consumption of curd be c & whey be w, then

Budget line : 5c + 3w = m

c/w = 1/2

w = 2c

putting this value into budgetl line

5c + 3*2 c = m

11c =m

c = m / 11